OpenStudy (anonymous):

A student needs to make a rectangular cardboard piece that is 12 inches long. The area of the cardboard piece should be between 60 square inches to 300 square inches. The function f(w) relates the area of the cardboard, in square inches, to the width w in inches. Which of the following shows a reasonable domain for f(w)? 2 ≤ w ≤ 5 5 ≤ w ≤ 25 25 ≤ w ≤ 60 60 ≤ w ≤ 300

OpenStudy (anonymous):

@johnweldon1993 can you help me

OpenStudy (johnweldon1993):

Alright...well we know Area = Length x Width And we want the total area to be between 60 and 300 sq in! So this means we have \[\large 60 \le \text{Length x Width} \ge 300\] We know length = 12 so we now have \[\large 60 \le 12 \times \text{Width} \ge 300\] Now in order to solve for width...we need to divide everything by 12 \[\large \frac{60}{12} \le \text{Width} \le \frac{300}{12}\] Simplify that out and we get what?

OpenStudy (anonymous):

B

OpenStudy (johnweldon1993):

Correct...did that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you help with 2 more